OpenStudy (anonymous):

solve for x |2x+7|=5

OpenStudy (anonymous):

make them two different equations

OpenStudy (anonymous):

2x+7=7 or 2x+7=-7

OpenStudy (anonymous):

can you get it from there?

OpenStudy (nurali):

|2x+7|=5 2x+7=5 or 2x+7=-5 2x=5-7 or 2x=-5-7 2x=-2 or 2x=-12 x=-1 or x=-6